OpenStudy (anonymous):

What is the solution to the equation? x = 1 x = 3 x = 5 x = 8

OpenStudy (anonymous):

\[3/2x+7=2^x\]

OpenStudy (anonymous):

@iGreen @SolomonZelman @ganeshie8

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{3}{2x}+7=2^x }\) like this /

OpenStudy (anonymous):

\[3/2x+7/2=2x\] Sorry i messed up

OpenStudy (solomonzelman):

how do you do these equations in class, what methods do you use may I know?

OpenStudy (solomonzelman):

oh, that is much better (lol)

OpenStudy (anonymous):

none...........check ur option

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{3}{2x}+7=2x }\)

OpenStudy (solomonzelman):

is my interpretation correct?

OpenStudy (anonymous):

It's 7/2 sorry

OpenStudy (anonymous):

It's \[3/2x+7/2=2^x\]

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{3}{2x} + \frac{7}{2}=2x }\) like this?

OpenStudy (solomonzelman):

you don't mean the x in the exponent, do you ?

OpenStudy (anonymous):

x is the exponent after the equal sign

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{3}{2x} + \frac{7}{2}=2^x }\)

OpenStudy (anonymous):

Yep

OpenStudy (solomonzelman):

I don't think there is an elementary way to solve this.

OpenStudy (solomonzelman):

You can plug in your options to check it.... or graph the function on a graphing calculator

OpenStudy (anonymous):

Ok, i'll try that

OpenStudy (anonymous):

it was x=3

OpenStudy (solomonzelman):

yes, it is x=3